OpenStudy (anonymous):

find the slope of the tangent to the curve of the intersection of the surface 4z= sqrt(25x^2+9y^2-225) and the plane y=1 at the point (3,1,3/4)

OpenStudy (anonymous):

If I can get answers to all these questions that would be great

myininaya (myininaya):

That seems to be a midterm exam.

myininaya (myininaya):

This doesn't seem very honest.

OpenStudy (anonymous):

myininaya this was an old midterm exam. I'm studying for my final. 5 questions from this midterm will be on the exam.

myininaya (myininaya):

Oh okay. I just wanted to make sure.

OpenStudy (anonymous):

@myininaya so can you help me out?

myininaya (myininaya):

I can probably help with one or two but I suggest asking one question at a time.

myininaya (myininaya):

The first question just wants you to find f_x and f_y

myininaya (myininaya):

\[\text{ to find } f_x \text{ we need to treat y like it is a constant }\] I will do this but I will leave finding f_y to you \[f=(\sin(4xy^2-y))^2 \\ f_x=2(\sin(4xy^2-y))^{2-1}\cos(4x^2-y) (8x-0)\] when it came to differentiating y I treated it like a constant

myininaya (myininaya):

but as you see I used chain rule like twice

myininaya (myininaya):

now the second order partial derivatives are the ones like f_xx ,f_yy, f_xy, and f_yx

myininaya (myininaya):

I will find one of those for you for the next function and leave the other 3 for you

myininaya (myininaya):

\[f=\cos(xy^2) \\ f_y=(x \cdot 2y)(-\sin(xy^2))=-2xysin(xy^2) \text{ treated x like constant } \\ \text{ \to find } f_{yx} \text{ we need to use product rule } \\ f_{yx}=(-2xy)_xsin(xy^2)+-2xy(\sin(xy^2))_x \\ =-2ysin(xy^2)-2xy \cdot y^2 \cos(xy^2)\]

myininaya (myininaya):

so notice I have to find first order partial to find second order partial

myininaya (myininaya):

are there any questions to what I have done?

myininaya (myininaya):

I could tell you how to do number 3 if you come back... but I will start you off if you do and I have left \[\text{ find } f_x \text{ and } f_y \text{ where } f(x,y)=z \\ \text{ then your normal vector would be } (f_x(3,1),f_y(3,1),-1) \\ \text{ your tangent plane should be } \\ f_x(3,1) \cdot (x-3)+f_y(3,1) \cdot (x-1)-1(z-\frac{3}{4}) =0 \\ \text{ if I'm not mistaken }\]

myininaya (myininaya):

4) see if you can get a different path to show a different limit from another path